OpenStudy (anonymous):

Can anybody explain the steps to the algebra that arrives at the derivative function to f[x] = x/e^(x^2) f'[x] = ???

OpenStudy (australopithecus):

are you using the definition of the derivative or are you allowed to use the rule set?

OpenStudy (logan13):

35

OpenStudy (anonymous):

I think at this point anything goes, I'm mainly just trying to understand how the f[x+h] - f[x] /h equation cancels out.. or if there is another alternative approach that gets me there. Its not actually for any test.. just my ability to use and apply it.

OpenStudy (logan13):

ya me 2

OpenStudy (australopithecus):

Here is a list of basic calc rules (calculus 1) If you want to solve the derivative quickly I would suggest learning the rules they are included on this pdf. I can help you use them and break it down for you if you want. Although if you want we can work through the definition of the derivative method, been a long time since I have dont it but shouldn't be too hard.

OpenStudy (anonymous):

okay I think that means the definition of the derivative for me

OpenStudy (anonymous):

ah thats awesome.. man that's a lot of rules.. geez, how do you guys memorize this?

OpenStudy (australopithecus):

Its actually not that hard just like learning algebra rules. YOu apply them a million times and you can just look at a function and see its derivative

OpenStudy (anonymous):

yeah, I can see how this would save a lot of fiddly algebra

OpenStudy (australopithecus):

So first you need to set up the expression do you know how to? Hint: for f(x+h) you just replace all x terms in the function with (x+h)

OpenStudy (anonymous):

yes 1 sec .. I can post how far I got

OpenStudy (australopithecus):

Doing the definition of this will probably be really dirty actually lol

OpenStudy (australopithecus):

but I will help you none the less :D

OpenStudy (anonymous):

\[\left( \frac{ x+h }{e^{(x+h)^{2}}} - \frac{x}{e^{x^{2}}} \right)* \frac{1}{h}\] \[\left( (x+h)e^{-(x+h)^2}-x e^{-x^{2}} \right) * \frac{1}{h}\] \[\frac{ (x+h)e^{-(x+h)^2}-x e^{-x^{2}} } {h} \] and a few other variations of that

OpenStudy (anonymous):

thanks I really really appreciate it.. this is going on 14 hours of head scratching for me now

OpenStudy (anonymous):

Im not sure if I want to keep a denominator or change it somehow, or if there is a good way to get rid of it.

hartnn (hartnn):

you've solved limits questions before? have you come across \(\Large \lim \limits_{x\to 0 } \dfrac{e^x -1}{x}\) ?

OpenStudy (anonymous):

theres a good question ... hmmm ... its been a while.. about a year, but I think I did that in my high school pre calc..

hartnn (hartnn):

good, remember L'Hopital's rule? that can be easily solved by it...

OpenStudy (anonymous):

I couldnt tell you off hand how to solve that..

OpenStudy (anonymous):

I can go look it up and refresh though

hartnn (hartnn):

by L'Hopital's rule \(\Large \lim \limits_{x\to 0 } \dfrac{e^x -1}{x} =\Large \lim \limits_{x\to 0 } \dfrac{\dfrac{d}{dx}[e^x -1]}{\dfrac{d}{dx}[x]} = \\ \Large \lim \limits_{x\to 0 } e^x = e^0 =1\) If you haven't learnt how to take the derivatove, you can always take \(\Large \lim \limits_{x\to 0 } \dfrac{e^x -1}{x} = 1\) as a standard result

hartnn (hartnn):

now with the algebra part, \(\Large (x+1) e^{-(x+1)^2 } - xe^{-x^2} \) expand (x+1)^2 if we look carefully, we can just factor out \(e^{-x^2}\) from that can you try that ?

OpenStudy (anonymous):

Yeah, I dont think I've covered this before.. but.. it looks like it makes some sense.. there are two opposing functions of some kind that are equal at extreme infitesimal values of x and therefore give the result one. And from this I guess we can simplify the equation. I'll go read up and fill in this blank in my understanding.

hartnn (hartnn):

x+h, not x+1 sorry :P

hartnn (hartnn):

i'll type the simplification by the time you try \(\Large (x+h) e^{-x^2 +2xh-h^2 } - xe^{-x^2}\) \(\Large (x+h) e^{-x^2} e^{2xh-h^2 } - xe^{-x^2}\) \(\Large e^{-x^2} [(x+h) e^{2xh-h^2 } - x]\)

OpenStudy (anonymous):

ahhh! you expanded the (x+h)^2 then used the e^(a + b) = e^a e^b rule to get to stage 2.. then factored out e^-x^2 .. jeez I was having a hell of time working out how to do that.. I got basic algebra holes

hartnn (hartnn):

thats right! :) now forget about e^(-x^2) that will be a constant and will be pulled out of the limit, as we are taking the limit with "h" as variable. \(\Large [(x+h) e^{2xh-h^2 } - x]= x(e^{2xh-h^2 }) + h (e^{2xh-h^2 } )-x \) now we have 3 terms and don't forget the h in the denominator

hartnn (hartnn):

the middle term is interesting \(\huge \dfrac{h(e^{2xh-h^2})}{h} = e^{2xh-h^2}\) since h->0 , you can directly plug in h =0 in that ^^

hartnn (hartnn):

from the other 2 terms, factor out the x \(\Large \dfrac{x(e^{2xh-h^2 }) -x}{h}= x[\dfrac{(e^{2xh-h^2 }) -1}{h} ]\) now doesn't that fraction look very similar to e^x-1 ;) I'll let you try further... will come later to check how much you could solve :)

hartnn (hartnn):

ask if you have any doubts in any explanation above

OpenStudy (anonymous):

I will. thanks, just digesting it..

OpenStudy (anonymous):

Im having trouble finishing this with a working function.. One of the problems I have is when I go back to this point.. I dont get the same result when I expand.. instead I get .. \[ \Large \left((x+h) e^{-(x+h)^2 } - xe^{-x^2}\right) * \frac{1}{h}\] \[ \Large \left((x+h) e^{-((x+h)(x+h)) } - xe^{-x^2}\right) * \frac{1}{h}\] \[ \Large \left( (x+h) e^{-x^2-2hx-h^2 } - xe^{-x^2}\right) * \frac{1}{h}\] \[ \Large \left( x e^{-x^2-2hx-h^2 } + h e^{-x^2-2hx-h^2 } - xe^{-x^2}\right) * \frac{1}{h}\] \[ \Large e^{-x^2} \left( x e^{-2hx-h^2 } + h e^{-2hx-h^2 } - x\right) * \frac{1}{h}\] And before I go any further may I confirm if I haven't got it wrong from the start?

OpenStudy (australopithecus):

Ok so I solved this problem I didnt use L'Hopital's rule. Not really following how the other guy is using it. This is a pretty weird problem, was kind of fun solving because the weird limit at the end. First off to solve this problem you need to know limit rules Here is a list of them http://math.oregonstate.edu/home/programs/undergrad/CalculusQuestStudyGuides/SandS/lHopital/limit_laws.html First off, your algebra looks fine. Step 1 cancel out h terms in your equation Step 2 you should notice that one of your terms does not have h in the numerator or denominator thus you can treat it as its own function and apply this rule called the addition law: \[\lim_{h \rightarrow a} (g(x) + f(x)) = \lim_{h \rightarrow a} g(x) + \lim_{h \rightarrow a} f(x)\] Also notice that you can treat x terms such as e^-x^2 and x as constants so thus you can simplify your problem even more using the rule called the constant law: \[\lim_{h \rightarrow a} c*f(x) = c*\lim_{h \rightarrow a} f(x)\] (in this case c would be x and e^-x^2 terms) Step 3 The term with no h in it you can solve its limit leaving you with only two terms, combine those terms factor out all the x terms you can. Then you will be left with a pretty weird limit, but it is solvable. Step 4 once you have the weird limit you are going to have to do some funky math tricks (completing the square and realizing and factoring out a term) to get it in the form of the definition of a derivative from there you can realize that the solution to the limit is the derivative of a specific exponential function. I just used derivative rules to solve the limit but you can use the definition. the link below goes over how to solve exponential functions using the definition of a derivative if you need help let me know: http://tutorial.math.lamar.edu/Classes/CalcI/DiffExpLogFcns.aspx If you need help with this final step let me know and if you are confused with any of the details I would be happy to walk you through it.

OpenStudy (australopithecus):

Give me a bit I need to look up a proof for the exponential rule using definition of a derivative

OpenStudy (australopithecus):

or you could just be content with d/dx e^(f(x)) = f'(x)e^(f(x))

OpenStudy (australopithecus):

Do you follow my steps though? they are kind of arbitrary but yeah

OpenStudy (anonymous):

thanks, I can see how I can get rid of one h easily.. in the middle term.. the other two, I'm not sure how easily they will go. \[\Large e^{-x^2} \left(\frac{ x e^{-2hx-h^2 }}{h} + \frac{ h e^{-2hx-h^2} }{h} - \frac{x}{h} \right) \] \[\Large e^{-x^2} \left(\frac{ x e^{-2hx-h^2 }}{h} + \frac{ e^{-2hx-h^2} }{1} - \frac{x}{h} \right) \] I feel like I need to do something about getting rid of those exponents already, do you have a trick for reducing xe^... any more ?

OpenStudy (australopithecus):

You need to use the limit notation while solving your problem

OpenStudy (australopithecus):

so right now, \[\lim_{h \rightarrow 0} e^{-x^2}(\frac{xe^{-2hx-h^2}}{h} + e^{-2hx-h^2} + \frac{x}{h})\]

OpenStudy (australopithecus):

now apply the rules I showed you above

OpenStudy (anonymous):

cool this is looking good, okay I might need a bit to read up on pauls notes there.. and get the gist of how the rules apply on this function as it is.. that might take me a bit.. then I have an appointment at 6 tonight., so it might be a few hours before I have this down.

OpenStudy (australopithecus):

Lets just try to get it in the right form first?

OpenStudy (anonymous):

feels like it's 1 minute from an ah ha! moment though

OpenStudy (australopithecus):

then we can try to tackle the limit

OpenStudy (australopithecus):

so yeah using proof to solve the limit is apparently very long and tricky according to some people who know what they are talking about. the link I posted wont be helpful in this case I dont think

OpenStudy (anonymous):

if I understand it right, this function will become several limit functions?

OpenStudy (australopithecus):

No you will split this into two functions

OpenStudy (anonymous):

and conceptually, x and e will become roughly the same very small value

OpenStudy (australopithecus):

you have 3 terms, you split them into two functions, then you solve one of the limits, then you combine the function with two terms and simplify

OpenStudy (australopithecus):

Can you apply the addition rule to your problem?

OpenStudy (anonymous):

...looking at that

OpenStudy (australopithecus):

If you want an example I can provide one?

OpenStudy (australopithecus):

For example \[\lim_{h \rightarrow 0} (x + \frac{x}{h})\] Applying addition rule, f(x) = x and g(x) = x/h so, \[\lim_{h \rightarrow 0} (x + \frac{x}{h}) = \lim_{h \rightarrow 0} x + \lim_{h \rightarrow 0} \frac{x}{h}\]

OpenStudy (australopithecus):

we can also apply the constant rule here by assuming x is a constant because we are taking the limit in terms of h so, \[\lim_{h \rightarrow 0} x + \lim_{h \rightarrow 0} \frac{x}{h} = x* \lim_{h \rightarrow 0} 1 + x*\lim_{h \rightarrow 0} \frac{1}{h} = x(\lim_{h \rightarrow 0} 1 + \lim_{h \rightarrow 0} \frac{1}{h})\] this isnt the best example because there is no limit for 1/h thus the addition rule is not valid but yeah just an example of how to do the manipulatons

OpenStudy (australopithecus):

or rather the limit of x/h does not exist

OpenStudy (anonymous):

ah ok is this where I should be next? \[\lim_{h \rightarrow 0} e^{-x^2} (\lim_{h \rightarrow 0} + \frac{xe^{-2hx-h^2}}{h} + \lim_{h \rightarrow 0} \frac{x}{h}+ \lim_{h \rightarrow 0} e^{-2hx-h^2} )\]

OpenStudy (australopithecus):

No sorry that is wrong, lets take a step back, you have three terms, and you want to treat all individual x terms as constants. so first off you want to designate an f(x) term and a g(x) term

OpenStudy (australopithecus):

look at the first part of my example

OpenStudy (australopithecus):

you want to group the terms with h in the denominator

OpenStudy (australopithecus):

using the rule, \[\frac{a}{b} + \frac{c}{b} = \frac{a+c}{b}\]

OpenStudy (australopithecus):

then you will have two terms

OpenStudy (australopithecus):

Then you can declare them as separate functions and apply the addition rule

OpenStudy (anonymous):

So first if I understand this right, then that term on the front can be ignored, because it does not have h in it, I think this is what you mean by constant? \[\lim_{h \rightarrow 0} \frac{xe^{-2hx-h^2 }+x}{h} + \lim_{h \rightarrow 0} e^{-2hx-h^2} \]

OpenStudy (anonymous):

or at least I am guessing for the time being, can be put on the side. and used later.

OpenStudy (australopithecus):

You should have e^(-x^2) multiplied by everything but yeah that is right

OpenStudy (anonymous):

ok, yes, that what I was thinking..

OpenStudy (australopithecus):

so, \[e^{-x^2}(\lim_{h \rightarrow 0} \frac{xe^{-2hx - h^2} +x}{h} + \lim_{h \rightarrow 0} e^{-2hx - h^2})\]

OpenStudy (australopithecus):

You can solve one of your terms leaving you with only one limit to solve

OpenStudy (australopithecus):

factor out x of the last remaining limit

OpenStudy (anonymous):

\[ e^{-x^2}(\lim_{h \rightarrow 0} \frac{x(e^{-2hx - h^2} +1)}{h} + \lim_{h \rightarrow 0} e^{-2hx - h^2}) \]

OpenStudy (australopithecus):

yeah now you can solve the limit of e^(-2hx - h^2)

OpenStudy (anonymous):

should I be thinking something along the lines of dropping that second part of that exponent.. (my signs might be off) but something like \[ e^{-x^2}(\lim_{h \rightarrow 0} \frac{x(e^{-2hx - h^2} +1)}{h} + \lim_{h \rightarrow 0} e^{-2hx}/e^{- h^2}) \]

OpenStudy (australopithecus):

for the second limit all you have to do is sub 0 into h

OpenStudy (australopithecus):

the last limit you need to manipulate it to look like the definition of a derivative

OpenStudy (anonymous):

-2x h - h*h = -2x *0 - 0*0 = 0 ? or do think of this as 2x and 0's cancel out in some way ?

OpenStudy (australopithecus):

no that is right you will have e^0 for the second limit

OpenStudy (australopithecus):

the reason we cant solve the first limit is because h is in the denominator and dividing by 0 is undefined

OpenStudy (anonymous):

oh cool, seemed too easy

OpenStudy (australopithecus):

so you have the first term of the derivative but to find the second you have to manipulate the last remaining limit to look like the definition of a derivative, hint complete the square in the exponential and factor

OpenStudy (anonymous):

\[ e^{-x^2}(\lim_{h \rightarrow 0} \frac{x(e^{-2hx - h^2} +1)}{h} + \lim_{h \rightarrow 0} e^{-2x* 0- 0 * 0}) \] \[ e^{-x^2}(\lim_{h \rightarrow 0} \frac{x(e^{-2hx - h^2} +1)}{h} + \lim_{h \rightarrow 0} e^{0}) \]

OpenStudy (australopithecus):

yeah so, you have the lim h->0 e^0 = 1

OpenStudy (australopithecus):

do you know how to complete the square?

OpenStudy (anonymous):

ah right I just spotted that ... \[ e^{-x^2}(\lim_{h \rightarrow 0} \frac{x(e^{-2hx - h^2} +1)}{h} + \lim_{h \rightarrow 0} 1) \]

OpenStudy (australopithecus):

you should remove the limit notation from the second one

OpenStudy (australopithecus):

since it is redundant

OpenStudy (anonymous):

yeah, bit rusty on it.. there's a couple of ways I recall. not sure if these terms divide nicely..

OpenStudy (australopithecus):

the last limit is a bit tricky

OpenStudy (australopithecus):

do you have any idea how to make it look like the definition of a deriative

OpenStudy (anonymous):

I have an idea of how it should look when done..

OpenStudy (australopithecus):

So the last limit you want it in the form of \[\frac{f(x+h) - f(x)}{h}\]

OpenStudy (australopithecus):

well when you are done you wll have a function in terms of x that is the derivative you are looking for

OpenStudy (anonymous):

\[f'[x] = ​(1 - 2*x^2)* e^{-x^2]} \]

OpenStudy (australopithecus):

huh?

OpenStudy (anonymous):

when I cheat and ask mathematica to give me the derivative of f[x] that's what it gives me.

OpenStudy (australopithecus):

ok so you have this right now \[e^{-x^2}(\lim_{h \rightarrow 0} \frac{e^{-2hx-h^2}-1}{h} + 1)\]

OpenStudy (australopithecus):

yeah that is the derivative notice when you expand you get e^(-x^2) which is showing up in the problem atm